OpenStudy (anonymous):

For the following system, use the second equation to make a substitution for x in the first equation. 3x + 2y = 7 x - y + 3 = 0 What is the resulting equation? 3x - y - 3 + 2y = 7 3(y - 3) + 2y = 7 3y - 3 + 2y = 7

OpenStudy (jackellyn):

x -y +3 = 0 x = y - 3 3(y-3) +2y = 7

OpenStudy (anonymous):

thx alot

OpenStudy (jackellyn):

no problem